# Structure-Accurate Medical Image Translation based on Dynamic Frequency Balance and Knowledge Guidance

### v1

@@ Appendix B Additional Experimental Results
As shown in Fig. 6, 7 and 8, our method consistently outperforms other methods for medical image translation on BraTs, ISLES, and SynthRAD datasets, demonstrating superior anatomical accuracy, and better quantitative metrics. Unlike other methods, our approach effectively maintains structural integrity, accurately locates lesion areas, and ensures realistic conversion between imaging modalities. Therefore, our method is robust and reliable in medical image translation and has great clinical application value.
Figure 6: Additional results on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of T1 to T2 and T1 to FLAIR translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
Figure 7: Additional results on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of FLAIR to DWI and FLAIR to T1 translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
@@
Figure 8: Additional results on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of CT to MR and MR to CT translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.

### v2

@@ Appendix B Additional Experimental Results
As shown in Fig. 5, 6, 7 and 8, our method consistently outperforms other methods for medical image translation on BraTs, ISLES, and SynthRAD datasets, demonstrating superior anatomical accuracy, and better quantitative metrics. Unlike other methods, our approach effectively maintains structural integrity, accurately locates lesion areas, and ensures realistic conversion between imaging modalities. Therefore, our method is robust and reliable in medical image translation and has great clinical application value.
@@
Figure 5: Comparison of approaches on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours). The first and third rows show the results of CT to MR and MR to CT translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively.
Figure 6: Additional results on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours). The first and third rows show the results of T1 to T2 and T1 to FLAIR translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
Figure 7: Additional results on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours). The first and third rows show the results of FLAIR to DWI and FLAIR to T1 translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
Figure 8: Additional results on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours). The first and third rows show the results of CT to MR and MR to CT translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.